# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ III-E The Update Step of $\underline{\mathbf{R}}$
III-F Box-Constrained Movement Mode for FAs
Although constraints (32) and (37b) are linear and compatible with quadratic optimization algorithms, they introduce $\frac{1}{2}M(M-1)$ and $\frac{1}{2}NK(N-1)$ inequalities, respectively. As the number of constraints grows proportionally to $M^{2}$, solving problem (12) becomes infeasible for large $M$. Moreover, the original movement mode presents challenges for praectical implmentation. Since all FAs share a common movable region, mechanical conflicts may arise, limiting the feasibility of the design in real-world applications [37, 38].
@@ III-F Box-Constrained Movement Mode for FAs
respectively. In problems (45) and (46), which consider the box-constrained movement mode, the total number of inequalities in the constraints grows linearly with the number of FAs. Specifically, the numbers are $M$ and $NK$, respectively.
@@
$p_{m}^{t}$ $=\min\Big{(}\max\left(p_{m}^{t,\star},p_{m}^{\min}\right),p_{m}^{\max}\Big{)},$ (47a)
$p_{kn}^{r}$ $=\min\Big{(}\max\left(p_{kn}^{r,\star},p_{kn}^{\min}\right),p_{kn}^{\max}\Big{)},$ (47b)
respectively, where $p$ denotes a spatial coordinate, and can be replaced by $x$, $y$, or $z$, depending on the dimension being optimized. Specifically, $p_{m}^{t,\star}$ and $p_{kn}^{r,\star}$ are the entries of the unconstrained optimal solutions $\mathbf{T}^{\star}$ and $\mathbf{R}_{k}^{\star}$, respectively.
The box-constrained movement mode restricts the feasible domain of the problem (12). Although this approach sacrifices some achievable WSR for reduced complexity, we will show in Section V-B that the degradation is negligible if the movable regions are sufficiently large. Based on the discussions above, we summarize the proposed BCA-based joint beamforming and antenna position optimization in Algorithm 1.
@@
1:$M$, $N$, $K$, $P_{\max}$, $\alpha_{k}$, $\mathbf{\Sigma}_{k}$, $L_{k}^{t}$, $L_{k}^{r}$, $\theta_{ki}^{t}$, $\phi_{ki}^{t}$, $\theta_{kj}^{r}$, $\phi_{kj}^{r}$.
@@
5:     Update each $\mathbf{W}_{k}$ by bisection search via (III-C).
6:     Update $\mathbf{T}$ using MM according to Section III-D.
@@
       Calculate $\delta^{\mathtt{Tx}}$ via (31).
@@
         Calculate $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ via (28).
@@
13:     Update $\mathbf{R}_{k}$ using MM according to Section III-E.
@@
21:$\underline{\mathbf{W}}$, $\mathbf{T}$, $\underline{\mathbf{R}}$.
IV Decentralized Implementation of the BCA-Based Algorithm
Although the proposed BCA-based algorithm in Section III can effectively solve the problem (12), its centralized implementation presents significant challenges as $M$ increases. The reasons are summarized as follows.
  • Storage cost: The CU requires substantial storage for CSI, beamforming matrices, and transmit FA positions, all of which have high dimensions.
  • Computational cost: Matrix multiplications and inversions become increasingly prohibitive as $M$ grows.
  • Interconnection cost: In multiuser systems, antennas must transmit each user’s CSI matrix to the CU, and the CU must send the position information matrix to the FAs. Transmitting such high-dimensional matrices requires significant overhead.
To address the challenges while maintaining similar performance, we propose a decentralized implementation of Algorithm 1 based on the DBP architecture.
IV-A DBP Architecture
Figure 2: DBP architecture for FA-assisted MU-MIMO systems.
The DBP architecture is illustrated in Fig. 2, where the black solid line represents the data signal, and the red dash-dot line indicates the control signal. The DBP architecture is implemented on the BS side and partitions the transmit FA array into $C$ clusters. Each cluster contains $M_{c}$ transmit FAs with $M=CM_{c}$, and is managed by a DU with dedicated RF circuitry, storage, and a baseband processor. The black solid line represents the data signal, while the red dash-dot line indicates the control signal. With the DBP architecture, problem (12) is solved cooperatively by the CU and DUs, requiring data exchange between them. Specifically, the DUs compute the beamforming matrices $\mathbf{W}_{k}$ and transmit FA positions $\mathbf{T}$, enabling both beamforming and FA position control at the DU.
Based on the DBP architecture, Algorithm 1 can be extended to its decentralized implementation that address the challenge above. Specifically, to alleviate the storage cost, any matrix of dimension $M$ is stored distributively at the DUs. The superscript $c$ denotes the submatrix stored across the $c$-th DU. Specifically, the variables requiring distributive storage include the channel matrix $\mathbf{H}_{k}$, the positions of transmit FAs $\mathbf{T}$, the FRMs for transmit FAs $\mathbf{G}_{k}$, the beamforming matrices $\mathbf{W}_{k}$, and the transpose of the first-order derivative of $f_{\rm Quad}$ w.r.t. $\mathbf{G}_{k}$, denoted as $\mathbf{D}_{k}^{\mathtt{Tx}}$.
To manage the computational cost, we must ensure that complexity at both the CU and DUs is independent of $M$. In the remainder of this section, we discuss the distributed computation of each set of variables in the proposed BCA-based method. Besides, the low interconnection cost can be naturally achieved with reduced storage cost, as the dimensions of all stored matrices are independent of $M$.
IV-B Decentralized Calculation of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$
@@
The values of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$, which are essential for evaluating system performance, are computed and stored at the CU. To compute them, we define $\tilde{\mathbf{G}}_{kj}\triangleq\mathbf{G}_{k}(\mathbf{T})\mathbf{W}_{j}$ and compute $\tilde{\mathbf{G}}_{kj}$ distributively. To calculate $\tilde{\mathbf{G}}_{kj}$, each DU first computes $\mathbf{G}_{k}^{c}(\overline{\mathbf{T}}^{c})\overline{\mathbf{W}}_{j}^{c}$ locally and then transmits the result to the CU. At the CU, $\tilde{\mathbf{G}}_{kj}$ is aggregated as
$\tilde{\mathbf{G}}_{kj}=\sum\nolimits_{c=1}^{C}\mathbf{G}_{k}^{c}\left(\overline{\mathbf{T}}^{c}\right)\overline{\mathbf{W}}_{j}^{c}.$ (48)
Once the CU obtains $\tilde{\mathbf{G}}_{kj}$, the values of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$ are obtained by
@@
$f_{\rm Quad}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}\right)=\sum_{k=1}^{K}\left(\alpha_{k}\log\det\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)-\alpha_{k}\mathtt{tr}\left(\mathbf{\Gamma}_{k}\right)\right.\\ \left.+\mathtt{tr}\left(\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\left(\sqrt{\alpha_{k}}\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\mathbf{F}_{k}\mathbf{\Phi}_{k}+\sqrt{\alpha_{k}}\mathbf{\Phi}_{k}^{\mathsf{H}}\mathbf{F}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kk}-\mathbf{\Phi}_{k}^{\mathsf{H}}\left(\sum_{j=1}^{K}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kk}\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}+\sigma_{k}^{2}\mathbf{I}\right)\mathbf{\Phi}_{k}\right)\right)\right).$ (51)
IV-C Decentralized Update of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
The values of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ are computed and stored at the CU. Additionally, they can be transmitted to and stored at the DUs. Computing $\mathbf{\Gamma}$ and $\mathbf{\Phi}$ also requires $\tilde{\mathbf{G}}_{kj}$, which has already been calculated by (48). With $\tilde{\mathbf{G}}_{kj}$, the remaining computations for updating $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ can be performed directly at the CU. The corresponding expressions are given by
@@ IV-C Decentralized Update of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
IV-D The Decentralized Update of $\underline{\mathbf{W}}$
$f_{\rm NonH}\left(\underline{\mathbf{W}},\underline{\mathbf{\Psi}}\right)=\sum_{k=1}^{K}\mathtt{tr}\left(-\eta\mathbf{W}_{k}^{\mathsf{H}}\mathbf{W}_{k}-2\Re\left(\mathbf{W}_{k}^{\mathsf{H}}\left(\sum_{j=1}^{K}{\mathbf{H}_{j}^{\mathsf{H}}\mathbf{\Phi}_{j}\left(\mathbf{I}+\mathbf{\Gamma}_{j}\right)\mathbf{\Phi}_{j}^{\mathsf{H}}\mathbf{H}_{j}}-\eta\mathbf{I}\right)\mathbf{\Psi}_{k}\right)\right.\\ -\left.\mathbf{\Psi}_{k}^{\mathsf{H}}\left(\eta\mathbf{I}-\sum_{j=1}^{K}{\mathbf{H}_{j}^{\mathsf{H}}\mathbf{\Phi}_{j}\left(\mathbf{I}+\mathbf{\Gamma}_{j}\right)\mathbf{\Phi}_{j}^{\mathsf{H}}\mathbf{H}_{j}}\right)\mathbf{\Psi}_{k}+\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\left(\sqrt{\alpha_{k}}\mathbf{W}_{k}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\mathbf{\Phi}_{k}+\sqrt{\alpha_{k}}\mathbf{\Phi}_{k}\mathbf{H}_{k}\mathbf{W}_{k}\right)\right)+\mathrm{const}.$ (58)
A major challenge in the decentralized update of $\underline{\mathbf{W}}$ is the matrix inversion in (III-C). Since matrix inversion of dimension $M$ cannot be computed distributively, it is necessary to avoid matrix inversion in (III-C) to achieve a decentralized update of $\underline{\mathbf{W}}$. Therefore, we first discuss an inverse-free update step of $\underline{\mathbf{W}}$, followed by the proposed decentralized implementation of the algorithm.

### v2

@@ IV-E Update Step of $\underline{\mathbf{R}}$
IV-F Box-Constrained Movement Mode for FAs
Although constraints (32) and (37b) are linear and compatible with quadratic optimization algorithms, they introduce $\frac{1}{2}M(M-1)$ and $\frac{1}{2}NK(N-1)$ inequalities, respectively. As the number of constraints grows proportionally to $M^{2}$, solving problem (11) becomes infeasible for large MM. The original movement mode also presents challenges for practical implementation. Since all FAs share a common movable region, mechanical conflicts may arise, limiting the feasibility of the design in real-world applications [48, 49]. Moreover, under the DBP architecture discussed in Section V, FAs from different clusters may violate the constraint (11e), potentially leading to physical collisions.
@@ IV-F Box-Constrained Movement Mode for FAs
respectively. In problems (45) and (46), which adopt the box-constrained movement mode, the total number of inequality constraints increases linearly with the number of FAs. Specifically, problems (45) and (46) contain $M$ and $NK$ inequality constraints, respectively.
@@
$p_{m}^{\mathtt{Tx}}$ $=\min\Big(\max\left(p_{m}^{t,\star},p_{m}^{\min}\right),p_{m}^{\max}\Big),$ (47a)
$p_{kn}^{\mathtt{Rx}}$ $=\min\Big(\max\left(p_{kn}^{r,\star},p_{kn}^{\min}\right),p_{kn}^{\max}\Big),$ (47b)
respectively, where $p$ denotes a spatial coordinate, and can be replaced by $x$, $y$, or $z$, depending on the dimension being optimized. Specifically, $p_{m}^{t,\star}$ and $p_{kn}^{r,\star}$ are the entries of the unconstrained optimal solutions $\mathbf{T}^{\star}$ and $\mathbf{R}_{k}^{\star}$, respectively. The box-constrained movement mode restricts the feasible domain of the problem (11). Although this approach sacrifices some achievable WSR for reduced complexity, we will show in Section VI-A that the degradation is negligible if the movable regions are sufficiently large.
Based on the discussions above, we summarize the proposed BCA-based joint beamforming and antenna position optimization in Algorithm 1. In step 2, the beamforming matrices are initialized as $\mathbf{W}_{k}=\sqrt{\frac{P_{\rm max}}{Kd}}\left[\mathbf{I}_{d},\mathbf{0}_{d\times(M-d)}\right]^{\mathsf{T}}$. Let $\rho$ denote the movable region of each antenna, normalized by $\lambda$, and the initial positions of transmit and receive FAs are uniform planar arrays with the spacing $\rho\lambda$. Notice that for the case of the decentralized implementation in Section V, all the parameters are initialized similarly.
@@
1:$M$, NN, $K$, $P_{\max}$, $\alpha_{k}$, $\mathbf{\Sigma}_{k}$, $L_{k}^{\mathtt{Tx}}$, $L_{k}^{\mathtt{Rx}}$, $\theta_{ki}^{\mathtt{Tx}}$, $\phi_{ki}^{\mathtt{Tx}}$, $\theta_{kj}^{\mathtt{Rx}}$, $\phi_{kj}^{\mathtt{Rx}}$.
@@
5:  Update each $\mathbf{W}_{k}$ by bisection search via (19).
6:  Update $\mathbf{T}$ using $M$ according to Section IV-D.
@@
    Calculate $\delta^{\mathtt{Tx}}$ via (IV-D).
@@
      Calculate $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ via (IV-D).
@@
13:  Update $\mathbf{R}_{k}$ using $M$ according to Section IV-E.
@@
[…]
IV-G Convergence Analysis
Since Algorithm 1 is a two-loop algorithm, its convergence is ensured by the convergence of the MM-based inner loop and the BCA-based outer loop. To show the convergence of the proposed MM-based inner loop, we only discuss the convergence of the MM for transmit FA position optimization $\mathbf{T}$ for brevity. The convergence of the proposed MM algorithm is ensured by the monotonic increase and the upper boundedness of the objective function $f_{\rm Quad}(\mathbf{T})$. We introduce the following lemmas to demonstrate the convergence of the proposed MM algorithm.
Lemma 2.
Let $\overline{\mathbf{T}}$ and $\mathbf{T}$ be the FA position matrix before and after an MM iteration, respectively. Then, the objective value $f_{\rm Quad}$ increases monotonically, i.e.,
$f_{\rm Quad}\left(\mathbf{T}\right)\geq f_{\rm Quad}\left(\overline{\mathbf{T}}\right),$ (48)
Proof:
According to the discussions in Section IV-D, the following inequalities hold:
where the first inequality follows from the surrogate function property (21), the second holds because $\mathbf{T}$ is the optimal solution to the problem (23), and the final equality results from the equality condition of the surrogate function at the expansion point $\overline{\mathbf{T}}$, given by (22). ∎
Lemma 3.
The objective function $f_{\rm Quad}$ has finite upper bound as long as the constraints (11b)–(11f) hold, i.e., $\exists R_{\max}\in\mathbb{R}_{+}$, such that
$f_{\rm Quad}(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}})\leq R_{\max}$ (50)
for all $\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}$ satisfying (11b)–(11f).
Proof:
Please refer to Appendix -C. ∎
Given the above two lemmas, we conclude that the objective function $f_{\rm Quad}$ is monotonically increasing and upper bounded within each MM iteration. According to [38, Proposition 4], the resulting solution $\mathbf{T}$ must be a stationary point, and the convergence of the proposed MM-based inner loop is thus guaranteed. The convergence of the BCA-based outer loop is a well-established result and is proved in [38, Section V-D]. Therefore, the overall convergence of Algorithm 1 is ensured.
V Decentralized Implementation of the BCA-Based Algorithm
Although the proposed BCA-based algorithm in Section IV can effectively solve the problem (11), this centralized implementation suffers from high computational costs as $M$ increases. The DBP architecture provides a promising solution to address the challenge by enabling DUs to solve small-scale subproblems in parallel. However, the advantages of the DBP architecture cannot be achieved without efficient decentralized optimization algorithms. To reduce the computational cost while maintaining similar performance, we propose a decentralized implementation of Algorithm 1 based on the DBP architecture.
V-A DBP Architecture
Figure 2: DBP architecture for FA-assisted MU-MIMO BS.
The DBP architecture for FA-assisted MU-MIMO system is illustrated in Fig. 2, where the black solid line represents the data signal, and the red dash-dotted line indicates the control signal. The DBP architecture is implemented on the BS side and partitions the transmit FA array into $C$ clusters. Each cluster contains $M_{c}$ transmit FAs with $M=CM_{c}$, and is managed by a DU equipped with dedicated RF circuitry, storage, and a baseband processor. With the DBP architecture, problem (11) is solved cooperatively by the centralized unit (CU) and DUs, requiring data exchange between them. Specifically, the DUs compute the beamforming matrices $\mathbf{W}_{k}$ and transmit FA positions $\mathbf{T}$, enabling both beamforming and FA position control at the DU. To alleviate storage and interconnection costs, any matrix of dimension $M$ is stored distributively at the DUs. The letter $c$ denotes the submatrix stored across the $c$-th DU. The variables requiring distributive storage include the channel matrix $\mathbf{H}_{k}$, the positions of transmit FAs $\mathbf{T}$, the FRMs for transmit FAs $\mathbf{G}_{k}$, the beamforming matrices $\mathbf{W}_{k}$, and the transpose of the first-order derivative of $f_{\rm Quad}$ w.r.t. $\mathbf{G}_{k}$, denoted as $\mathbf{D}_{k}^{\mathtt{Tx}}$. Since each DU has access only to the positions of the transmit FAs it optimizes, FAs from different DUs may violate the constraint (11e), potentially resulting in physical collisions. To address this issue, we adopt the box-constraint movement mode introduced in Section IV-F, which assigns independent and non-overlapping movable regions to each FA. To manage the computational cost, we must ensure that the complexity at both the CU and DUs grows as a function of per-cluster antenna-count $M_{c}$, instead of MM. [^6] In practical systems, the number of DUs $C$ is typically fixed by the hardware design. The complexity at the CU and DUs is still proportional to $M$. However, as long as there is a sufficient number of DUs $C$, the complexity at the CU and DUs is no longer dominated by $M$.
Figure 3: The decentralized calculation of $\sum_{c=1}^{C}(\mathbf{A}^{c})^{\mathsf{H}}\mathbf{B}^{c}$.
Decentralized computation methods form the foundation of the proposed BCA-based decentralized algorithm. Among these methods, decentralized matrix multiplication is particularly critical and will be discussed in detail later. Given its pivotal role in our proposed algorithm, we first formalize this operation along with its implementation before presenting the algorithmic details. Suppose matrices $\mathbf{A}^{c}\in\mathbb{C}^{M_{c}\times a}$ and $\mathbf{B}^{c}\in\mathbb{C}^{M_{c}\times b}$ are stored at the $c$-th DU. We define the function $\mathsf{Mul}(\cdot,\cdot)$ as
$\mathsf{Mul}(\mathbf{A}^{c},\mathbf{B}^{c})=\sum\nolimits_{c=1}^{C}(\mathbf{A}^{c})^{\mathsf{H}}\mathbf{B}^{c},$ (51)
whose result has dimensions independent of MM and can be efficiently stored and transmitted between the CU and DUs. The decentralized computation of $\mathsf{Mul}(\mathbf{A}^{c},\mathbf{B}^{c})$ is illustrated in Fig. 3. Specifically, each DU first computes $(\mathbf{A}^{c})^{\mathsf{H}}\mathbf{B}^{c}$ locally and then transmits the result to the CU, where the final sum is aggregated.
V-B Decentralized Calculation of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$
@@
The values of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$, which are essential for evaluating system performance, are computed and stored at the CU. To compute them, we define $\tilde{\mathbf{G}}_{kj}\triangleq\mathsf{Mul}\Big(\big(\mathbf{G}^{c}_{k}(\overline{\mathbf{T}}^{c})\big)^{\mathsf{H}},\overline{\mathbf{W}}_{j}^{c}\Big)$. Once the CU obtains $\tilde{\mathbf{G}}_{kj}$ as illustrated in Fig. 3, the values of $\mathbf{M}_{k}$, $R$, and $f_{\rm Quad}$ are calculated directly at the CU:
@@
$f_{\rm Quad}\left(\underline{\mathbf{W}},\mathbf{T},\underline{\mathbf{R}},\underline{\mathbf{\Gamma}},\underline{\mathbf{\Phi}}\right)=\sum_{k=1}^{K}\left(\alpha_{k}\log\det\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)-\alpha_{k}\mathtt{tr}\left(\mathbf{\Gamma}_{k}\right)\right.\\ +\mathtt{tr}\Big[\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\Big(\sqrt{\alpha_{k}}\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\mathbf{F}_{k}\mathbf{\Phi}_{k}+\sqrt{\alpha_{k}}\mathbf{\Phi}_{k}^{\mathsf{H}}\mathbf{F}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kk}-\mathbf{\Phi}_{k}^{\mathsf{H}}\Big(\sum_{j=1}^{K}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kk}\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}+\sigma_{k}^{2}\mathbf{I}\Big)\mathbf{\Phi}_{k}\Big)\Big]\Big).$ (54)
V-C Decentralized Update of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
The values of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ are computed and stored at the CU. Additionally, they can be transmitted to and stored at the DUs. With $\tilde{\mathbf{G}}_{kj}$, the remaining computations for updating $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$ can be performed directly at the CU:
@@ V-C Decentralized Update of $\underline{\mathbf{\Gamma}}$ and $\underline{\mathbf{\Phi}}$
V-D Decentralized Update of $\underline{\mathbf{W}}$
$f_{\rm NonH}\left(\underline{\mathbf{W}},\underline{\mathbf{\Psi}}\right)=\sum_{k=1}^{K}\mathtt{tr}\Big[-\eta\mathbf{W}_{k}^{\mathsf{H}}\mathbf{W}_{k}-2\Re\Big(\mathbf{W}_{k}^{\mathsf{H}}\Big(\sum_{j=1}^{K}{\mathbf{H}_{j}^{\mathsf{H}}\mathbf{\Phi}_{j}\left(\mathbf{I}+\mathbf{\Gamma}_{j}\right)\mathbf{\Phi}_{j}^{\mathsf{H}}\mathbf{H}_{j}}-\eta\mathbf{I}\Big)\mathbf{\Psi}_{k}\Big)\\ -\mathbf{\Psi}_{k}^{\mathsf{H}}\Big(\eta\mathbf{I}-\sum_{j=1}^{K}{\mathbf{H}_{j}^{\mathsf{H}}\mathbf{\Phi}_{j}\left(\mathbf{I}+\mathbf{\Gamma}_{j}\right)\mathbf{\Phi}_{j}^{\mathsf{H}}\mathbf{H}_{j}}\Big)\mathbf{\Psi}_{k}+\left(\mathbf{I}+\mathbf{\Gamma}_{k}\right)\left(\sqrt{\alpha_{k}}\mathbf{W}_{k}^{\mathsf{H}}\mathbf{H}_{k}^{\mathsf{H}}\mathbf{\Phi}_{k}+\sqrt{\alpha_{k}}\mathbf{\Phi}_{k}\mathbf{H}_{k}\mathbf{W}_{k}\right)\Big]+\mathrm{const}.$ (61)
A major challenge in the decentralized update of $\underline{\mathbf{W}}$ is the matrix inversion in (19). Since matrix inversion of dimension $M$ cannot be computed distributively, it is necessary to avoid matrix inversion in (19) to achieve a decentralized update of $\underline{\mathbf{W}}$. Therefore, we first discuss an inverse-free update step of $\underline{\mathbf{W}}$, followed by the proposed decentralized implementation of the algorithm.